2sin(pi/4x)=1/2. Find all solutions
Answer by lwsshak3(11628)

2sin(pi/4x)=1/2. Find all solutions
***
2sin(pi/4x)=1/2
sin(pi/4x)=1/4
inverse sin(1/4)≈0.2527
(π/4)x=0.2527
x=.2527/(π/4)≈0.3217